OpenStudy (anonymous):

A jar contains 8 green, 5 red, 4 blue, and 3 black marbles.once a marble is selected, it is not replaced. find each probability. 1) p(two green marbles) 2) p(two black marbles) 3) p(a red marble and then a blue marble) 4) p(a black marble then a green marble) 5) p(two marbles that are not red) 6) p(two marbles that are neither blue nor green) 7) a family has two children, both of whom are boys. what is the probability that the next child will be born a boy?

OpenStudy (anonymous):

Hi! First understand the sample space - all possibilities.

OpenStudy (anonymous):

Sample space equals 8 green + 5 red + 4 blue + 3 black = 20. This will always be the denominator of the probability. remember, probability is: Desired Outcome/All Outcome

OpenStudy (anonymous):

two black marbles means 1) first is black and 2) second is black given that the first is black multiply these two probabilities together first is black \(\frac{3}{20}\) second is black given that the first is black \(\frac{2}{19}\) your answer \[\frac{3}{20}\times \frac{2}{19}\]

OpenStudy (anonymous):

not to be argumentative, but 20 is not always the denominator

OpenStudy (anonymous):

Yes, I agree. But this is assuming we are only picking one marble. I understand, without replacement, the denominator decreases :)

OpenStudy (anonymous):

thats the part im confused at

OpenStudy (anonymous):

For part 1: Two green marbles means a green marble on the first try AND a green marble on the second try.

OpenStudy (anonymous):

@helpmeplease123123 is it clear how i solved number two? because the others are similar

OpenStudy (anonymous):

Since there is an "and" involved, we use the multiplication principle of probability. P(2 green) = P(1 green 1st try)*P(1 green 2nd try)

OpenStudy (anonymous):

im lost still :(

OpenStudy (anonymous):

first is black \(\frac{3}{20}\) because there are 3 black and 20 in total now given that the first one is black, there are only 2 black ones left, and only 19 total so the probability that the second one is black given that the first one is black is \(\frac{2}{19}\)

OpenStudy (anonymous):

so for 1) is 1 over 19

OpenStudy (anonymous):

For part 1: Two green marbles means a green marble on the first try AND a green marble on the second try. Since there is an "and" involved, we use the multiplication principle of probability. P(2 green) = P(1 green 1st try)*P(1 green 2nd try) There are 8 green marbles. First try probability: P(1 green 1st try) = Desired/Total = (8 green)/(20 total) = 2/5 Now, there are only 19 marbles left since 1 green is gone.

OpenStudy (anonymous):

so the probability that both are black is the product of those two probabilities \[\frac{3}{20}\times \frac{2}{19}\]

OpenStudy (anonymous):

soo 1 over 19 ?

OpenStudy (anonymous):

p(1 green 2nd try) = desired/total = 7 green/19 total = 7/19 Final probability equals (2/5)*(7/19)=15/85=3/17

OpenStudy (anonymous):

oooo thanks

OpenStudy (anonymous):

This is for number 1. The rest are similar. Take a look at mine and satellite's solution for #1 and #2.

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

i still need help with the rest